Could you please tell me if I am correct on this equation?
4-v/6-v=2/v-6. Am I correct in thinking there is no solution?
Thanks for all your help!!

Answer by stanbon(75887)

4-v/6-v=2/v-6
---
Rearrange:
(v-4)/(v-6) = 2/(v-6)
----
Multiply thru by (v-6) to get:
v-4 = 2
v = 6
------
But v cannot be 6 because that would
make the orignal fractions undefined.
--------------------
You are correct; no solution.
==============
Cheers,
Stan H.